# references.bib

### v1

@@ 1.1 Mean-field neural networks and two-timescale learning
  • At fixed $\lambda>0$, the teacher-student assumption that $Y=\Phi\star\bar{\nu}$ is not restrictive since one can always replace $Y$ by its orthogonal projection on the set $\{\Phi\star\nu,\,\nu\in\mathcal{M}(\Omega)\}$, thereby only modifying $\mathcal{L}^{\lambda}$ by subtracting a constant term. However, this assumption becomes crucial in the limit $\lambda\to 0^{+}$ to ensure the feasibility of the the optimization problem in eq. 10.
  • The injectivity assumption on $\Phi\star$ ensures uniqueness of the reference measure $\bar{\nu}$. In the limit where $\lambda\to 0^{+}$, this allows to rewrite the $\mathcal{L}^{0}$ only in terms of a divergence between $\bar{\nu}$ and $\mu$ (eq. 17). In the case $\lambda>0$, $\mathcal{L}^{\lambda}$ is an infimal convolution between this divergence and a kernel discrepency (eq. 18) and the injectivity assumption ensures this discrepency is a distance on the space of measures (Lemma A.1). It will be useful in section 5 to prove convergence of Wasserstein gradient flows of $\mathcal{L}^{\lambda}$ to solutions of ultra-fast diffusion equations. In the case of a two-layer neural network when the feature map is of the form $\phi((w,b),x)=\sigma(w^{\top}x+b)$ the injectivity assumption is satisfied as soon as $\sigma$ is not a polynomial and the data distribution has full support on $\mathbb{R}^{d}$ ([sun2019random, Thm. III.4]).
@@ 1.2 Contributions and related works
This paper studies the convergence of gradient methods for the training of a mean-field model of neural network in the two-timescale regime. Precisely, we study the dynamic of the feature distribution $\mu\in\mathcal{P}(\Omega)$ when trained with gradient flow for the minimization of the reduced risk $\mathcal{L}^{\lambda}$, for $\lambda\geq 0$. In the teacher-student scenario defined by Assumption 1, we establish guarantees for the convergence of $\mu$ towards the teacher feature distribution $\bar{\mu}$:
  • In the case $\lambda=0$, we show in section 4 that this dynamic corresponds to an ultra-fast diffusion equation. Relying on the work of [iacobelli2019weighted], this allows to state convergence towards the teacher feature distribution $\bar{\mu}$ (Theorem 4.3), with a linear convergence rate.
  • At fixed $\lambda>0$, we establish in Theorem 5.1 convergence of $\mu$ towards the teacher feature distribution $\bar{\mu}$ with an algebraic rate.
  • Finally, in the limit $\lambda\to 0^{+}$, we show under regularity assumptions that, in the two-timescale regime, the dynamic of the feature distribution $\mu$ converges locally uniformly in time to the solution of the ultra-fast diffusion equation with weights $\bar{\mu}$ (Theorem 5.2).
@@
Several works have studied the convergence of gradient based methods for the training of neural networks models similar to eq. 1 with the mean-field scaling $\frac{1}{M}$. In [bach2021gradient], the authors show that for two layer neural networks with a homogeneous activation, if gradient flow on the weights distribution converges then it converges towards a global minimizer of the risk. In [rotskoff2019global], a similar result is shown for a modification of the gradient flow dynamic where a supplementary “birth-death” term is added.
Several works have also analyzed the convergence of noisy gradient descent, or Langevin dynamic, for the training of mean-field models of two layer neural networks [chizat2022mean, mei2019mean, nitanda2022convex, hu2021mean, suzuki2023feature]. Thanks to the addition of an entropic regularization, these work provide a convergence rate for the sampling of an invariant weight distribution.
@@
While two-timescales learning strategies have a broad range of application in the fields of stochastic approximation and optimization [borkar1997stochastic, borkar2008stochastic], there has been a recent interest in these methods for the training of neural networks [marion2023leveraging, berthier2024learning, wang2024mean, bietti2023learning, takakura2024mean]. Specifically, [berthier2024learning] studies the training of a two-layer neural network and exhibits a separation of timescales and different learning phases whose respective sizes depends the timescale parameter $\gamma$. In [marion2023leveraging], the authors study two-timescale gradient descent for a simple model of $1$-dimensional neural network and show that the teacher network is recovered as soon as both the number of neurons of the student and the timescale parameter are sufficiently large. In [bietti2023learning], a multi-index regression problem is considered. Relying on the assumption of high dimensional gaussian data, the authors consider a linear layer composed with a non-parameteric model whose projection can be computed in the Hermite basis. They show this VarPro algorithm results in a saddle-to-saddle dynamic on the linear layer and establish guarantees for the recovery of the teacher model.
@@
Wasserstein gradient flow of statistical distances
Under our Assumption 1, eq. 18 shows $\mathcal{L}^{\lambda}$ is an infimal convolution of statistical divergences between the feature distribution $\mu$ and the teacher $\bar{\nu}$, interpolating between the $\chi^{2}$-divergence $\chi^{2}(\bar{\nu}|\mu)$ (or more generally a $f$-divergence $\mathrm{D}_{f}(\bar{\nu}|\mu)$) when $\lambda\to 0^{+}$ and a (squared) kernel discrepancy $\mathrm{MMD}(\bar{\nu},\mu)^{2}$ when $\lambda\to\infty$. In the case $\lambda\to\infty$, gradient flows of $\mathrm{MMD}$-discrepancies and applications to sampling were studied in several works [arbel2019maximum, sejdinovic2013equivalence, hertrich2023wasserstein, hertrich2023generative, hertrich2024wasserstein, boufadene2023global]. Those flows are known to get trap in local minima but discrepancies associated to non-smooth kernel have been observed to behave better in terms of convergence [hertrich2023generative, hertrich2024wasserstein]. In the case of the coulomb kernel, [boufadene2023global] proves that the discrepancy loss admits no spurious local minima and that the discrepancy flow converges towards the target measure under regularity assumptions.
@@ 1.3 Mathematical preliminaries and notations
In the following, $\Omega$ will either be the $n$-dimensional Torus or a closed bounded convex domain of $\mathbb{R}^{n}$, for some $n\geq 1$. We note $\mathcal{M}(\Omega)$ the set of finite Borel measure over $\Omega$ and $\mathcal{P}(\Omega)$ is the subset of $\mathcal{M}(\Omega)$ consisting of probability measures. In particular, we will denote by $\pi\in\mathcal{P}(\Omega)$ the uniform distribution over $\Omega$. For a measure $\nu\in\mathcal{M}(\Omega)$, $|\nu|$ is its total-variation measure and $\|\nu\|_{\mathrm{TV}}$ is the total variation of $\nu$. For $p\in[1,+\infty)$, we note $\mathcal{W}_{p}$ the Wasserstein-$p$ distance defined for two probability measures $\mu,\mu^{\prime}\in\mathcal{P}(\Omega)$ by:
$\mathcal{W}_{p}(\mu,\mu^{\prime})\coloneqq\min_{\gamma\in\Gamma(\mu,\mu^{\prime})}\left(\int_{\Omega\times\Omega}\|\omega-\omega^{\prime}\|^{2}\mathrm{d}\gamma(\omega,\omega^{\prime})\right)^{1/p}$
@@
For an integer $k\geq 0$ and $s\in(0,1]$, we note $\mathcal{C}^{k,s}(\Omega)$ (or just $\mathcal{C}^{k,s}$ when no ambiguity) the Hölder space of $k$-times continuously differentiable functions over $\Omega$ with $s$-Hölder $k^{\text{th}}$-derivative. We note .$\|.\|_{\mathcal{C}^{k,s}}$ the Hölder norm on $\mathcal{C}^{k,s}(\Omega)$. For a probability measure $\rho\in\mathcal{P}(\mathbb{R}^{d})$ and $p\in[1,+\infty]$, we note $L^{p}(\rho,\mathcal{C}^{k,s})$ the space of measurable functions $\phi:\Omega\times\mathbb{R}^{d}\to\mathbb{R}$ s.t. $\phi(.,x)\in\mathcal{C}^{k,s}(\Omega)$ for $\mathrm{d}\rho$-a.e. $x\in\Omega$ and $\|\phi\|_{L^{p}(\rho,\mathcal{C}^{k,s})}\coloneqq\left(\int_{\mathbb{R}^{d}}\|\phi(.,x)\|^{p}_{\mathcal{C}^{k,s}}\mathrm{d}\rho(x)\right)^{1/p}<+\infty$. We will often use that if $\phi\in L^{2}(\rho,\mathcal{C}^{k,s})$ and $\alpha\in L^{2}(\rho)$ then the Böchner integral $\int_{\mathbb{R}^{d}}\phi(.,x)\alpha(x)\mathrm{d}\rho(x)$ is in $\mathcal{C}^{k,s}$ with:
@@ 2 Reduced risk associated to the VarPro algorithm
We study in this work a VarPro algorithm or two-timescale limit of gradient descent for the training of neural networks. This strategy amounts to perform gradient descent on the reduced risk defined as the result of a partial minimization on a regularized version of the risk.
@@ 2.1 Primal formulation of the reduced risk
Whereas regularizing the risk with the euclidean square norm of the weights is a popular practice, the variable projection procedure can be used with other kind of regularization. Generally, for a convex function $f:\mathbb{R}\to\mathbb{R}$ and a regularization strength $\lambda>0$ we consider for $\mu\in\mathcal{P}(\Omega)$ and $u\in L^{1}(\mu)$:
$\mathcal{R}^{\lambda}_{f}(\mu,u)\coloneqq\frac{1}{2}\left\|\Phi_{\mu}\cdot u-Y\right\|^{2}_{L^{2}(\rho)}+\lambda\int_{\Omega}f(u)\mathrm{d}\mu,$
@@
In the following, we always assume that $\phi\in L^{2}(\rho,\mathcal{C}^{0}(\Omega))$ (Assumption 3). This in particular implies that, for any $\mu\in\mathcal{P}(\Omega)$, the map $\Phi_{\mu}:L^{1}(\mu)\to L^{2}(\rho)$ is weakly continuous. We also consider the following assumptions on the regularization function:
@@
By Lemma 2.1, this is sufficient to ensure the existence of a unique minimizer $u^{\lambda}_{f}[\mu]\in\operatorname*{arg\,min}\mathcal{R}^{\lambda}_{f}(\mu,u)$ when $\lambda>0$ and $u^{0}_{f}[\mu]\in\operatorname*{arg\,min}_{\Phi_{\mu}\cdot u=Y}\int_{\Omega}f(u)\mathrm{d}\mu$ when $\lambda=0$. Of particular interest in this work and more precisely in section 4 will be the case where $f(t)=|t|^{r}/(r-1)$ for some $r>1$. In this case we will note the corresponding reduced risk $\mathcal{L}^{\lambda}_{r}$. In particular, for $r=2$ we recover the “$L^{2}$-regularized” reduced risk defined in eq. 4 and eq. 5.
@@
is strictly convex, weakly lower-semicontinuous and has weakly compact sublevel sets. In particular, eq. 9 (and eq. 10 if feasible) admits a unique minimizer $u^{\lambda}_{f}[\mu]$.
@@
Clearly $\mathcal{I}_{f}$ is strictly convex. Weak lower-semicontinuity is a classical consequence of the fact that $\mathcal{I}_{f}$ is convex and strongly lower-semicontinuous (using Fatou’s lemma), hence its epigraph is convex and strongly closed and hence also weakly closed. For weak compacity of sublevel sets, if $(u_{n})_{n\geq 0}$ is a sequence s.t. $\int_{\Omega}f(u_{n})\mathrm{d}\mu\leq C$ for every $n\geq 0$, then using that $f$ has super-linear growth, for every $\varepsilon>0$ there exists a $T\geq 0$ s.t. $|t|\leq\varepsilon f(t)$ for every $|t|\geq T$ and for every $n\geq 0$:
@@ 2.2 Partial minimization on the space of measures
The reduced risk can also be obtain as the result of partial minimization of a convex functional over the space of measures. Whereas we have previously separated the role of the outer weights $u$ and of the feature distribution $\mu$ in eq. 2, our neural network model can equivalently be seen as a convolution of a measure on the feature space $\Omega$.
@@
In the case $\lambda>0$, $\mathcal{L}^{\lambda}_{f}$ can be interpreted as the infimal convolution between a Maximum Mean Discrepancy (MMD) and the divergence $\mathrm{D}_{f}$. Indeed, naturally associated to the data distribution $\rho\in\mathcal{P}(\mathbb{R}^{d})$ and to the feature map $\phi$ is a structure of Reproducing Kernel Hilbert Space (RKHS) of functions that we will note $\mathcal{H}$ with a kernel $\kappa:\Omega\times\Omega\to\mathbb{R}$ defined by:
@@
We refer to appendix A for results of the theory of RKHSs we will use in this paper. Importantly, it follows from the definition of $\kappa$ and $\mathcal{H}$ that, under Assumption 1, the data attachment term in eq. 16 can be interpreted as a kernel distance between $\nu$ and $\bar{\nu}$. By eq. 43 we have $\|\Phi\star(\nu-\bar{\nu})\|_{L^{2}(\rho)}=\mathrm{MMD}_{\kappa}(\nu,\bar{\nu})$ where $\mathrm{MMD}_{\kappa}$ is the Maximum Mean Discrepancy (MMD) with kernel $\kappa$ [muandet2017kernel, gretton2012kernel]. For $\lambda>0$, the functional $\mathcal{L}^{\lambda}_{f}$ can then be expressed for every $\mu\in\mathcal{P}(\Omega)$ as:

### v2

@@ 1.1 Mean-field neural networks and two-timescale learning
  • At fixed $\lambda>0$, the teacher-student assumption that $Y=\Phi\star\bar{\nu}$ is not restrictive since one can always replace $Y$ by its orthogonal projection on the set $\{\Phi\star\nu,\,\nu\in\mathcal{M}(\Omega)\}$, thereby only modifying $\mathcal{L}^{\lambda}$ by subtracting a constant term. However, this assumption becomes crucial in the limit $\lambda\to 0^{+}$ to ensure the feasibility of the optimization problem in Eq. 18.
  • The injectivity assumption on $\Phi\star$ ensures uniqueness of the reference measure $\bar{\nu}$. In the limit where $\lambda\to 0^{+}$, this allows rewriting $\mathcal{L}^{0}$ only in terms of a divergence between $\bar{\nu}$ and $\mu$ (Eq. 25). In the case $\lambda>0$, $\mathcal{L}^{\lambda}$ is an infimal convolution between this divergence and a kernel discrepancy (Eq. 26) and the injectivity assumption ensures this discrepancy is a distance on the space of measures (Lemma A.1). It will be useful in Section 5 to prove convergence of Wasserstein gradient flows of $\mathcal{L}^{\lambda}$ to solutions of the ultra-fast diffusion equation. In the case of a two-layer perceptron, the feature map is of the form $\phi((w,b),x)=\sigma(w^{\top}x+b)$ and the injectivity assumption is satisfied as soon as $\sigma$ is not a polynomial and the data distribution has full support on $\mathbb{R}^{d}$ ([sun2019random, Thm. III.4]).
@@ 1.2 Contributions and related works
This paper studies the convergence of the VarPro algorithm — or two-timescale regime of gradient descent — for the training of mean-field models of neural networks. Precisely, we study the dynamic of the feature distribution $\mu\in\mathcal{P}(\Omega)$ when trained with gradient flow for the minimization of the reduced risk $\mathcal{L}^{\lambda}$, for $\lambda\geq 0$. In the teacher-student scenario defined by Assumption 1, we establish guarantees for the convergence of $\mu$ towards the teacher feature distribution $\bar{\mu}$:
  • In the case $\lambda=0$, we show in Section 4 that the training dynamic corresponds to an ultra-fast diffusion equation. Relying on the work of [iacobelli2019weighted], this allows stating convergence towards the teacher feature distribution $\bar{\mu}$ (Theorem 3), with a linear convergence rate.
  • At fixed $\lambda>0$, we establish in Theorem 4 convergence of $\mu$ towards the teacher feature distribution $\bar{\mu}$ with an algebraic rate.
  • In the limit $\lambda\to 0^{+}$, we show that, under regularity assumptions, the dynamic of the feature distribution $\mu$ converges locally uniformly in time to the solution of the ultra-fast diffusion equation with weights $\bar{\mu}$ (Theorem 5).
  • Finally, we show in Section 6 that numerical results on low-dimensional learning problems with synthetic data are well-aligned with our theory. Overall, these experiments indicate that, when the regularization is sufficiently low, the VarPro dynamic indeed enters an “ultra-fast diffusion regime” where the student feature distribution converges to the teacher’s at a linear rate. We also show with experiments on CIFAR10 that the VarPro algorithm can be adapted to the training of more complex architectures such as ResNets and achieves generalization on supervised learning problems with large datasets.
@@
Several works have studied the convergence of gradient based methods for the training of neural network models similar to Eq. 1 with the mean-field scaling $\frac{1}{M}$. In [bach2021gradient], the authors show that, for two layer neural networks with a homogeneous activation, if gradient flow on the weights distribution converges then it converges towards a global minimizer of the risk. In [rotskoff2019global], a similar result is shown for a modification of the gradient flow dynamic where a supplementary “birth-death” term is added.
Several works have also analyzed the convergence of noisy gradient descent, or Langevin dynamic, for the training of mean-field models of two layer neural networks [chizat2022mean, mei2019mean, nitanda2022convex, hu2021mean, suzuki2023feature]. Thanks to the addition of an entropic regularization term, these works provide a convergence rate for the sampling of an invariant weight distribution.
@@
While two-timescale learning strategies have a broad range of applications in the fields of stochastic approximation and optimization [borkar1997stochastic, borkar2008stochastic], there has been a recent interest in these methods for the training of neural networks [marion2023leveraging, berthier2024learning, wang2024mean, bietti2023learning, takakura2024mean]. Specifically, [berthier2024learning] studies the training of two-layer neural networks and exhibits a separation of timescales and different learning phases whose respective sizes depend on the timescale parameter $\eta$. In [marion2023leveraging], the authors study two-timescale gradient descent for a simple model of $1$-dimensional neural network and show that the teacher network is recovered as soon as both the number of neurons of the student and the timescale parameter are sufficiently large. In [bietti2023learning], a multi-index regression problem is considered. Relying on the assumption of high dimensional Gaussian data, the authors consider a linear layer composed with a nonparameteric model whose projection can be computed in the Hermite basis. They show this VarPro algorithm results in a saddle-to-saddle dynamic on the linear layer and establish guarantees for the recovery of the teacher model.
@@
Wasserstein gradient flows of statistical distances
Under our Assumption 1, Eq. 26 shows $\mathcal{L}^{\lambda}$ is an infimal convolution of statistical divergences between the feature distribution $\mu$ and the teacher $\bar{\nu}$, interpolating between the $\chi^{2}$-divergence $\chi^{2}(\bar{\nu}|\mu)$ — or more generally a $f$-divergence $\mathrm{D}_{f}(\bar{\nu}|\mu)$ — when $\lambda\to 0^{+}$ and a (squared) kernel discrepancy $\mathrm{MMD}(\bar{\nu},\mu)^{2}$ when $\lambda\to\infty$. In the case $\lambda\to\infty$, gradient flows of $\mathrm{MMD}$-discrepancies and applications to sampling were studied in several works [arbel2019maximum, sejdinovic2013equivalence, hertrich2023wasserstein, hertrich2023generative, hertrich2024wasserstein, boufadene2023global]. Those flows are known to get trapped in local minima but discrepancies associated to non-smooth kernels have been observed to behave better in terms of convergence [hertrich2023generative, hertrich2024wasserstein]. In the case of the coulomb kernel, [boufadene2023global] proves that the discrepancy loss admits no spurious local minima and that the discrepancy flow converges towards the target measure under regularity assumptions.
@@ 1.3 Mathematical preliminaries and notations
In the following, $\Omega$ will either be the $n$-dimensional torus or a closed bounded convex domain of $\mathbb{R}^{n}$, for some $n\geq 1$. We denote by $\mathcal{M}(\Omega)$ the set of finite Borel measures over $\Omega$ and by $\mathcal{P}(\Omega)$ the subset of $\mathcal{M}(\Omega)$ consisting of probability measures. We will denote by $\pi\in\mathcal{P}(\Omega)$ the uniform distribution over $\Omega$. For a measure $\nu\in\mathcal{M}(\Omega)$, $|\nu|$ is its total variation measure and $\|\nu\|_{\mathrm{TV}}$ is the total variation of $\nu$. For $p\in[1,+\infty)$, we denote by $\mathcal{W}_{p}$ the Wasserstein-$p$ distance defined for two probability measures $\mu,\mu^{\prime}\in\mathcal{P}(\Omega)$ by:
$\mathcal{W}_{p}(\mu,\mu^{\prime})\coloneqq\min_{\gamma\in\Gamma(\mu,\mu^{\prime})}\left(\int_{\Omega\times\Omega}\|\omega-\omega^{\prime}\|^{2}\mathrm{d}\gamma(\omega,\omega^{\prime})\right)^{1/p}\,,$
@@
For an integer $k\geq 0$ and for $s\in(0,1]$, we denote by $\mathcal{C}^{k,s}(\Omega)$ (or just $\mathcal{C}^{k,s}$) the Hölder space of $k$-times continuously differentiable real-valued functions over $\Omega$ with $s$-Hölder $k^{\text{th}}$-derivative. We denote by .$\|.\|_{\mathcal{C}^{k,s}}$ the Hölder norm on $\mathcal{C}^{k,s}(\Omega)$. For a probability measure $\rho\in\mathcal{P}(\mathbb{R}^{d})$ and $p\in[1,+\infty]$, we denote by $L^{p}(\rho,\mathcal{C}^{k,s})$ the space of measurable functions $\phi:\Omega\times\mathbb{R}^{d}\to\mathbb{R}$ s.t. $\phi(.,x)\in\mathcal{C}^{k,s}(\Omega)$ for $\mathrm{d}\rho$-a.e. $x\in\Omega$ and $\|\phi\|_{L^{p}(\rho,\mathcal{C}^{k,s})}\coloneqq\left(\int_{\mathbb{R}^{d}}\|\phi(.,x)\|^{p}_{\mathcal{C}^{k,s}}\mathrm{d}\rho(x)\right)^{1/p}<+\infty$. We will often use that if $\phi\in L^{2}(\rho,\mathcal{C}^{k,s})$ and $\alpha\in L^{2}(\rho)$ then the Bochner integral $\int_{\mathbb{R}^{d}}\phi(.,x)\alpha(x)\mathrm{d}\rho(x)$ is in $\mathcal{C}^{k,s}$ with:
@@ 2 Reduced risk associated to the VarPro algorithm
We study in this work a VarPro algorithm or two-timescale regime of gradient descent for the training of neural networks. This strategy amounts to performing gradient descent on the reduced risk defined as the result of a partial minimization on a regularized version of the risk.
@@ 2.1 Primal formulation of the reduced risk
Whereas regularizing the risk with the Euclidean square norm of the weights is a popular practice, the variable projection procedure can be used with other kinds of regularization. Generally, for a convex function $f:\mathbb{R}\to\mathbb{R}$ and a regularization strength $\lambda>0$ we consider for $\mu\in\mathcal{P}(\Omega)$ and $u\in L^{1}(\mu)$:
$\mathcal{R}^{\lambda}_{f}(\mu,u)\coloneqq\frac{1}{2}\left\|F_{\mu,u}-Y\right\|^{2}_{L^{2}(\rho)}+\lambda\int_{\Omega}f(u)\mathrm{d}\mu=\frac{1}{2}\left\|\Phi_{\mu}\cdot u-Y\right\|^{2}_{L^{2}(\rho)}+\lambda\int_{\Omega}f(u)\mathrm{d}\mu,$ (16)
@@
In the following, we always assume that $\phi\in L^{2}(\rho,\mathcal{C}^{0}(\Omega))$ (Assumption 3). This in particular implies that, for any $\mu\in\mathcal{P}(\Omega)$, the map $\Phi_{\mu}:L^{1}(\mu)\to L^{2}(\rho)$ is weakly continuous. We also consider the following assumption on the regularization function:
@@
By Lemma 2.1, this is sufficient to ensure the existence of a unique minimizer $u^{\lambda}_{f}[\mu]\in\operatorname*{arg\,min}\mathcal{R}^{\lambda}_{f}(\mu,u)$, when $\lambda>0$, and $u^{0}_{f}[\mu]\in\operatorname*{arg\,min}_{\Phi_{\mu}\cdot u=Y}\int_{\Omega}f(u)\mathrm{d}\mu$, when $\lambda=0$. Of particular interest in this work and more precisely in Section 4 is the case where $f(t)=|t|^{r}/(r-1)$ for some $r>1$. In this case we denote the corresponding reduced risk by $\mathcal{L}^{\lambda}_{r}$. In particular, for $r=2$ we recover the “$L^{2}$-regularized” reduced risk defined in Eq. 12 and Eq. 13.
@@
is strictly convex, weakly lower semicontinuous and has weakly compact sublevel sets. In particular, Eq. 17 (and Eq. 18 if feasible) admits a unique minimizer $u^{\lambda}_{f}[\mu]$.
@@
Clearly $\mathcal{I}_{f}$ is strictly convex. Weak lower semicontinuity is a classical consequence of the fact that $\mathcal{I}_{f}$ is convex and strongly lower semicontinuous (using Fatou’s lemma), hence its epigraph is convex and strongly closed and hence also weakly closed. For weak compacity of sublevel sets, if $(u_{n})_{n\geq 0}$ is a sequence s.t. $\int_{\Omega}f(u_{n})\mathrm{d}\mu\leq C$ for every $n\geq 0$, then using that $f$ has super-linear growth, for every $\varepsilon>0$ there exists a $T\geq 0$ s.t. $|t|\leq\varepsilon f(t)$ for every $|t|\geq T$ and for every $n\geq 0$:
@@ 2.2 Partial minimization on the space of measures
The reduced risk can also be obtained as the result of partial minimization of a convex functional over the space of measures. Whereas we have previously separated the role of the outer weights $u$ and of the feature distribution $\mu$ in Eq. 2, our neural network model can equivalently be seen as a linear operator acting on the space $\mathcal{M}(\Omega)$ of finite measures on $\Omega$.
@@
In the case $\lambda>0$, $\mathcal{L}^{\lambda}_{f}$ can be interpreted as the infimal convolution between a Maximum Mean Discrepancy (MMD) and the divergence $\mathrm{D}_{f}$. Indeed, naturally associated to the data distribution $\rho\in\mathcal{P}(\mathbb{R}^{d})$ and to the feature map $\phi$ is a structure of Reproducing Kernel Hilbert Space (RKHS) of functions on $\Omega$. We refer to appendix A for results on the theory of RKHSs we use in this chapter. The RKHS $\mathcal{H}$ is defined in Eq. 61, and corresponds to the kernel $\kappa:\Omega\times\Omega\to\mathbb{R}$ defined by:
@@
It then follows from the definition of $\kappa$ and $\mathcal{H}$ that, under Assumption 1, the data attachment term in Eq. 24 can be interpreted as a kernel distance between $\nu$ and $\bar{\nu}$. By Eq. 63 we have $\|\Phi\star(\nu-\bar{\nu})\|_{L^{2}(\rho)}=\mathrm{MMD}_{\kappa}(\nu,\bar{\nu})$ where $\mathrm{MMD}_{\kappa}$ is the Maximum Mean Discrepancy (MMD) with kernel $\kappa$ [muandet2017kernel, gretton2012kernel]. For $\lambda>0$, the functional $\mathcal{L}^{\lambda}_{f}$ can then be expressed for every $\mu\in\mathcal{P}(\Omega)$ as: